The following reaction is important to produce a mixture known as ‘water gas’:

C (s)+H2O (g) ⇌ CO (g)+ H2 (g)

Suppose 0.100 mol of H2O and solid carbon, together with 7.16 atm. of CO and 8.17 atm. of H2, are placed in a 1.00 dm3 reactor at 800°C. The value of Kp is 14.1 under these conditions.

Use the ideal gas law to determine the partial pressures of all the species at the start, and then determine the concentrations of all the species at equilibrium.



1
Expert's answer
2020-08-23T07:43:42-0400

p0(H2O)=nRT/V=8,8 atm. If the change in partial pressure is x, then Kp=(7,16+x)*(8,17+x)/(8,8-x) From this equation x=2,08. So c(H2O)=p/RT=0,076 M c(CO)=0,105M c(H2)=0,116 M